OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, 1) and (2, 5)?

OpenStudy (anonymous):

\[\frac{y_2-y_1}{x_2-x_1}\] Say \[(x_1,y_1) \:\;\;is \:\;\; (3,1)\] and \[(x_2,y_2) \:\;\;is \:\;\; (2,5)\] Note: it does not matter which one is the first or second, they will both end up being the same slope, just be consistent throughout the problem. So: \[\frac{y_2-y_1}{x_2-x_1}=\frac{5-1}{2-3}=\frac{4}{-1}=-4\]